OpenStudy (90890908967777878888):

exprees 4 8\9 as an inproper fraction

OpenStudy (tanya123):

4 9/8?

OpenStudy (anonymous):

40/9 :)

OpenStudy (anonymous):

40/9 :)

OpenStudy (90890908967777878888):

tjhx

jimthompson5910 (jim_thompson5910):

\[\Large \color{red}{4} \frac{\color{green}{8}}{\color{blue}{9}} = \frac{\color{red}{4}*\color{blue}{9}+\color{green}{8}}{\color{blue}{9}} = \frac{44}{9}\]

OpenStudy (90890908967777878888):

what answer is it

OpenStudy (anonymous):

@jim_thompson5910 hmm. I think your right. lol